A vehicle travelling along a street skidded on an icy patch and hit a roadside barrier the time interval of the impact was 0.955 seconds during which the barrier exerted a force of 505 and what is the magnitude of the change in momentum from the vehicle


1
Expert's answer
2021-11-14T17:14:03-0500

"\Delta p=F \Delta t,"

"\Delta p=505\cdot 0.955=482~\frac{kg\cdot m}s."